If z = 1.67,  H0  can be rejected at ______________.

α = .01 and α = .05

α = .05 and α = .10     

α = .10, but not at α = .05

α = .05, but not at α = .10
Answer #1

Given that, z test statistic is, z = 1.67

Case I) For one-tailed test (right-tailed test)

p-value= P(Z > 1.67)=1 - P(Z < 1.67)=1 - 0.9525 = 0.0475

=> p-value = 0.0475

Since, p-value = 0.0475 is less than α = 0.05 and α = 0.10 but greater than α = 0.01.

So, we reject H0 at α = 0.05 and α = 0.10 but not reject at α = 0.01

Case 2) For two-tailed test

p-value= 2*P(Z > 1.67)= 2 * 0.0475 = 0.095

=> p-value = 0.095

Since, p-value = 0.095 is greater than α = 0.01 and α = 0.05 but less than α = 0.10

So, we reject H0 at α = 0.10 but not reject at α = 0.01 and α = 0.05